h. The following sample observations were randomly selected.

X 4 5 3 6 10

Y 4 6 5 7 7

Determine the regression equation.



1
Expert's answer
2021-11-16T16:04:03-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2\\
 \hline
 & 4 & 4 & 16 & 16 & 16\\
 & 5 & 6 & 30 & 25 & 36\\
 & 3 & 5 & 15.8 & 9 & 25\\
 & 6 & 7 & 42 & 36 & 49\\
 & 10 & 7 & 70 & 100 & 49\\
 Sum= & 28 & 29 & 173 & 186 & 175\\
\end{array}""\sum_iX_i=28, \sum_iY_i=29"

"\sum_iX_iY_i=173,\sum_iX_i=186, \sum_iY_i=175""\bar{X}=\dfrac{1}{n}\sum_iX_i=\dfrac{28}{5}=5.6"

"\bar{Y}=\dfrac{1}{n}\sum_iY_i=\dfrac{29}{5}=5.8"

"SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum_iX_i)^2=100-\dfrac{(28)^2}{5}"

"=29.2"

"SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum_iY_i)^2=175-\dfrac{(29)^2}{5}"

"=6.8"

"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum_iX_i)(\sum_iY_i)"

"=173-\dfrac{28(29)}{5}=10.6"

"m=slope=\dfrac{SS_{XY}}{SS_{XX}}"

"=\dfrac{10.6}{29.2}=0.3630137"

"n=\bar{Y}-m\bar{X}"

"=5.8-0.3630137(5.6)"

"=3.7671233"

Therefore, we find that the regression equation is:


"Y=3.7671233+0.363013X"



Correlation coefficient:


"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}}\sqrt{SS_{YY}}}=\dfrac{10.6}{\sqrt{29.2}\sqrt{6.8}}"

"=0.752246"


Strong positive correlation.